ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 46A22 Danet, Rodica Mihaela, Technical University of Civil Engineering of Bucharest, Romania: The extension of some linear operators. See [1] for an introductory story in this subject. New results concerning the extension of a linear operator T : G ⟶ Y (having possibly another properties) will be presented. G is a certain linear subspace of a certain ordered (topological) vector space X. T must be extended to X preserving all its properties. Firstly, Y will be a complete vector lattice, as usually, and then Y will be a certain ordered vector space with the Cantor property. Reference: [1] R. M. Danet, Some technics for the existence and the extension of linear operators, Publishing House of the Romanian Academy, Bucharest, 1993 (in Romanian), 212p.